# Graph Quantum Magic Squares and Free Spectrahedra

### v1

@@ 2.1 Quantum Magic Squares
For $n=1,2,3$ one has $\mathcal{C}^{(n)}=\mathcal{P}^{(n)}$. In particular, every $3\times 3$ quantum permutation matrix is commuting (see, e.g., [WB23, Lemma 2.5]).
@@
Let $\{R_{s}\}_{s\in\mathbb{N}}$ with $R_{s}\subseteq\operatorname{Mat}_{n}\bigl(\operatorname{Her}_{s}(\mathbb{C})\bigr)$, and set $R:=\bigcup_{s\in\mathbb{N}}R_{s}$. We say that $R$ is matrix convex if, for any $s,t\in\mathbb{N}$, any $A^{(1)},\dots,A^{(k)}\in R_{s}$, and any $V_{1},\dots,V_{k}\in\operatorname{Mat}_{s,t}(\mathbb{C})$ with
@@ 2.2 Free spectrahedra and linear matrix inequalities
That is, $\mathcal{D}_{A}(s)$ is the set of all $g$-tuples of $s\times s$ real Hermitian matrices $X$ such that the evaluation $L_{A}(X)$ is positive semidefinite.
@@
The free spectrahedron associated with $A$ is the set
@@
We now briefly recall the notion of dilations and Arveson extreme points, following [DDN20] and [EH19].
Given $X\in\mathcal{D}_{A}(s)$, an $(s+\ell)\times(s+\ell)$ block matrix
$Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in\mathcal{D}_{A}(s+\ell)$
is called a dilation of $X$. The dilation is trivial if $\beta=0$, in which case $Y$ is simply $X$ with an additional direct summand.
@@
A matrix $X\in\mathcal{D}_{A}(s)$ is an Arveson extreme point if every dilation of $X$ inside $\mathcal{D}_{A}$ is trivial; that is, whenever
$\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in\mathcal{D}_{A}(s+\ell),\qquad\ell\geq 1,$
one must have $\beta=0$.
We conclude this preliminary section by recalling some facts that will be relevant later. In her thesis, [VB22] provides a sketch of the proof that the set of quantum magic squares forms a compact free spectrahedron.
In [EH19] it was shown that every compact free spectrahedron is generated, under matrix convex combinations, by its Arveson extreme points, which we denote by $\operatorname{ArvesonExt}(\mathcal{D})$. For convenience, we recall the result here:
@@
Let $\mathcal{D}$ be a compact free spectrahedron. Then
$\mathcal{D}=\mathrm{mconv}\bigl(\operatorname{ArvesonExt}(\mathcal{D})\bigr),$
that is, the matrix convex hull of the Arveson extreme points of $\mathcal{D}$ coincides with the entire spectrahedron.
@@ 2.3 Quantum Automorphism of Graphs
Crucially, for some graphs, this algebra is non-commutative, implying the existence of "genuine quantum symmetries". A graph $\Gamma$ is said to have no quantum symmetry if the algebra is commutative, which implies $C(\mathrm{Aut}^{+}(\Gamma))=C(\mathrm{Aut}(\Gamma))$.
@@ 3.3 Quantum Magic Squares as Free Spectrahedra
The set $\mathcal{M}^{(n)}_{s}$ is a compact free spectrahedron. Specifically, it can be described as:
$\mathcal{M}^{(n)}_{s}\;=\;\Bigl\{\,Y=(Y_{ij})_{i,j=1}^{n}\in\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))\mid I_{n^{2}}\otimes I_{s}+\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}\otimes Y_{ij}\ \succeq\ 0\,\Bigr\},$
where the matrices $A_{ij}$ are the coefficient matrices arising from the affine parametrization.
@@ 3.4 Graph Quantum Magic Squares as Free Spectrahedra
Let $\Gamma$ be a $k$-regular graph and $s\geq 1$. Then the set of graph quantum magic squares $\mathcal{M}^{(\Gamma)}_{s}$ admits the monic LMI representation
$\mathcal{M}^{(\Gamma)}_{s}\;=\;\bigl\{\,Y\in\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))\mid I_{n^{2}}\otimes I_{s}+\sum_{(i^{\prime},j^{\prime})\in\mathcal{I}_{\mathrm{ind}}}B_{i^{\prime}j^{\prime}}\otimes Y_{i^{\prime}j^{\prime}}\ \succeq\ 0\,\bigr\}.$
where $\mathcal{I}_{\mathrm{ind}}$ represents the set of indices of the independent variables.
In particular, $\mathcal{M}^{(\Gamma)}=\bigcup_{s\geq 1}\mathcal{M}^{(\Gamma)}_{s}$ is a compact free spectrahedron, hence a matrix convex set.
@@ 3.5 GQMS and Arveson extreme points
Having established in Theorem 3.9 that $\mathcal{M}^{(\Gamma)}$ is a compact free spectrahedron, we may now apply 2.10 to conclude that
@@
In the non-graph case, [DDN20] showed that every quantum permutation matrix is an Arveson extreme point of the free spectrahedron $\mathcal{M}^{(n)}$. Since graph quantum permutation matrices form a subclass of quantum permutation matrices subject only to additional linear constraints the same argument used in [DDN20] also applies in the graph setting.
@@
Let $U\in\mathcal{P}^{(\Gamma)}$. Then $U\in\mathcal{P}^{(n)}$, so by [DDN20, Corollary 20] $U$ is an Arveson extreme point of $\mathcal{M}^{(n)}$.
@@ 3.5.1 Concluding Remarks on the Free-Spectrahedral Setting
The set $\mathcal{M}^{(\Gamma)}$ is therefore a compact free spectrahedron, defined by finitely many linear matrix inequalities with scalar coefficients. By the general theory of free spectrahedra (see Theorem 2.10), its matrix convex hull is generated by its Arveson extreme points.
Our Corollary 3.11 shows that every graph quantum permutation matrix belongs to $\operatorname{ArvesonExt}(\mathcal{M}^{(\Gamma)})$. Consequently, in analogy with the non-graph case [DDN20, Corollary 20], the extremal structure of $\mathcal{M}^{(\Gamma)}$ is strictly richer than that of $\mathcal{P}^{(\Gamma)}$ whenever the Birkhoff–von Neumann property fails for $\Gamma$.
@@ 4 Future Directions
For the non-graph case, [DDN20] [EH19] show that
@@ Acknowledgements
I warmly thank Makoto Yamashita for his continuous guidance, for many helpful discussions, and for his supervision throughout this project. I am also grateful to Inga Valentiner-Branth, and Tim Netzer for their clarifications regarding some technical aspects of their work, and for their kind replies to my questions.

### v2

@@ 2.1 Quantum Magic Squares
For $n=1,2,3$ one has $\mathcal{C}^{(n)}=\mathcal{P}^{(n)}$. In particular, every $3\times 3$ quantum permutation matrix is commuting (see, e.g., [WB23]*Lemma 2.5).
@@
Let $\{R_{s}\}_{s\in\mathbb{N}}$ with $R_{s}\subseteq\operatorname{Mat}_{n}\bigl(\operatorname{Her}_{s}(\mathbb{C})\bigr)$, and set $R:=\bigcup_{s\in\mathbb{N}}R_{s}$. We say that $R$ is matrix convex if, for any $s,t\in\mathbb{N}$, any $A^{(1)},\dots,A^{(k)}\in R_{s_{i}}$, and any $V_{1},\dots,V_{k}\in\operatorname{Mat}_{s_{i},t}(\mathbb{C})$ with
@@ 2.2 Free spectrahedra and linear matrix inequalities
That is, $\mathcal{D}_{A}(s)$ consists of all $g$-tuples of $s\times s$ Hermitian matrices $X$ such that the evaluation $L_{A}(X)$ is positive semidefinite.
@@
The free spectrahedron associated with $A$ is the graded set
@@
We now recall the notions of dilations and Arveson extreme points, following [DDN20] and [EH19].
Given $X\in\mathcal{D}_{A}(s)$, a block matrix
$Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}\in\mathcal{D}_{A}(s+\ell),\qquad\ell\geq 1,$
is called a dilation of $X$. The dilation is trivial if $\beta=0$, in which case $Y$ is a direct sum of $X$ with another element of $\mathcal{D}_{A}$.
@@
Let $\mathcal{D}=\bigcup_{s\geq 1}\mathcal{D}(s)$ be a matrix convex set. An element $X\in\mathcal{D}(s)$ is an Arveson extreme point of $\mathcal{D}$ if it admits no nontrivial dilation inside $\mathcal{D}$. That is, whenever
$Y\in\mathcal{D}(s+\ell)\quad\text{and}\quad Y=\begin{pmatrix}X&\beta\\ \beta^{*}&\gamma\end{pmatrix}$,
for some $\ell\geq 1$, one must have $\beta=0$.
Remark 2.10.
In the literature (see e.g. [EH19]), an Arveson extreme point is usually required to be irreducible, meaning that the underlying tuple admits no nontrivial common reducing subspace. Without this additional requirement, Definition 2.9 characterizes direct sums of Arveson extreme points. Since all extremal objects considered in this paper are irreducible, this distinction will not play a role in what follows.
Remark 2.11 (Arveson extreme points and spanning).
Let $\mathcal{D}$ be a compact free spectrahedron which is closed under complex conjugation. In particular, this hypothesis holds whenever the defining pencil has real coefficient matrices (since then $X\in\mathcal{D}(s)$ implies $\overline{X}\in\mathcal{D}(s)$). This will be the case for the spectrahedra arising from quantum magic squares and from the graph constraints considered later, since the corresponding coefficient matrices can be chosen real. By [EH19]*Theorem 1.3, such a set satisfies
$\mathcal{D}=\operatorname{mconv}\bigl(\operatorname{AbsExt}(\mathcal{D})\bigr),$
where $\operatorname{AbsExt}(\mathcal{D})$ denotes the set of absolute extreme points.
Since every absolute extreme point is, by definition, an Arveson extreme point, one has
$\operatorname{AbsExt}(\mathcal{D})\subseteq\operatorname{ArvesonExt}(\mathcal{D}).$
As $\mathcal{D}$ is matrix convex, this implies
$\mathcal{D}=\operatorname{mconv}\bigl(\operatorname{ArvesonExt}(\mathcal{D})\bigr).$
@@
Let $\mathcal{D}$ be a compact free spectrahedron which is closed under complex conjugation. Then
$\mathcal{D}=\operatorname{mconv}\bigl(\operatorname{ArvesonExt}(\mathcal{D})\bigr).$
@@ 2.3 Quantum Automorphism of Graphs
Crucially, for some graphs, this algebra is non-commutative, implying the existence of genuine quantum symmetries. A graph $\Gamma$ is said to have no quantum symmetry if the algebra is commutative, which implies $C(\mathrm{Aut}^{+}(\Gamma))=C(\mathrm{Aut}(\Gamma))$.
@@ 3.3 Quantum Magic Squares as Free Spectrahedra
The set $\mathcal{M}^{(n)}$ of quantum magic squares is affinely isomorphic to a compact free spectrahedron. More precisely, after an affine translation, it can be described as
$\widehat{\mathcal{M}}^{(n)}_{s}=\Bigl\{\,Y=(Y_{ij})_{i,j=1}^{n}\in\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))\,\Big|\,I_{n^{2}}\otimes I_{s}+\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}\otimes Y_{ij}\succeq 0\Bigr\},$
where the matrices $A_{ij}$ arise from the affine parametrization of the magic relations.
A closely related viewpoint is given in [MR4680468], where quantum magic squares are identified with the maximal matrix convex set associated to the (centered) Birkhoff polytope. More precisely, in their notation one has $\mathcal{M}^{(n)}_{s}=(B_{N})_{\max}(s)$ via an explicit description in terms of linear inequalities. Our approach here provides an explicit monic LMI (after an affine translation) tailored to the magic relations, which we will later adapt to incorporate graph-commutation constraints.
@@ 3.4 Graph Quantum Magic Squares as Free Spectrahedra
Let $\Gamma$ be a $k$-regular graph and let $s\geq 1$. There exists an affine translation
$\widehat{\mathcal{M}}^{(\Gamma)}_{s}=\mathcal{M}^{(\Gamma)}_{s}-X^{\circ}$
such that the translated set admits a monic LMI representation of the form
$\widehat{\mathcal{M}}^{(\Gamma)}_{s}=\bigl\{\,Y\in\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))\mid I_{n^{2}}\otimes I_{s}+\sum_{(i^{\prime},j^{\prime})\in\mathcal{I}_{\mathrm{ind}}}B_{i^{\prime}j^{\prime}}\otimes Y_{i^{\prime}j^{\prime}}\succeq 0\,\bigr\},$
where $\mathcal{I}_{\mathrm{ind}}$ denotes the set of indices corresponding to the independent variables in the affine parametrization.
In particular, the graded set
$\mathcal{M}^{(\Gamma)}=\bigcup_{s\geq 1}\mathcal{M}^{(\Gamma)}_{s}$
is affinely isomorphic to a compact free spectrahedron. Hence $\mathcal{M}^{(\Gamma)}$ is a compact matrix convex set.
@@ 3.5 GQMS and Arveson extreme points
Having established in Theorem 3.9 that $\mathcal{M}^{(\Gamma)}$ is affinely isomorphic to a compact free spectrahedron, we may now apply Theorem 2.12 to conclude that
@@
In the non-graph case, [DDN20] showed that every quantum permutation matrix is an Arveson extreme point of the free spectrahedron $\mathcal{M}^{(n)}$. Since graph quantum permutation matrices form a subclass of quantum permutation matrices subject only to additional linear constraints, the same argument used in [DDN20] also applies in the graph setting.
@@
Let $U\in\mathcal{P}^{(\Gamma)}$. Then $U\in\mathcal{P}^{(n)}$, so by [DDN20]*Corollary 20 $U$ is an Arveson extreme point of $\mathcal{M}^{(n)}$.
@@ 3.5.1 Concluding Remarks on the Free-Spectrahedral Setting
The set $\mathcal{M}^{(\Gamma)}$ is therefore affinely isomorphic to a compact free spectrahedron, defined by finitely many linear matrix inequalities with scalar coefficients. By the general theory of free spectrahedra (see Theorem 2.12), its matrix convex hull is generated by its Arveson extreme points.
Our corollary 3.11 shows that every graph quantum permutation matrix belongs to the set $\operatorname{ArvesonExt}(\mathcal{M}^{(\Gamma)})$. Consequently, in analogy with the non-graph case [DDN20]*Corollary 20, the extremal structure of $\mathcal{M}^{(\Gamma)}$ is strictly richer than that of $\mathcal{P}^{(\Gamma)}$ whenever the Birkhoff–von Neumann property fails for $\Gamma$.
@@ 4 Future Directions
For the non-graph case, [DDN20, EH19] show that
@@ Acknowledgements
I warmly thank Makoto Yamashita for his continuous guidance, for many helpful discussions, and for his supervision throughout this project. I am also grateful to Inga Valentiner-Branth, and Tim Netzer for their clarifications regarding some technical aspects of their work, and for their kind replies to my questions. I would also like to thank Tea Štrekelj for carefully reading an earlier version of the manuscript and for her helpful comments, in particular on aspects related to matrix convexity and free spectrahedra.